If the coefficient of x3 in the expression of (1+x)n is 20, find n and n is in power.

Grade:11

1 Answers

Arun
6 years ago
Dear Priyansh
 
General term = nCr x^r
Now
x³ coefficient = 20
r = 3
 
Now
nC3 = 20
Hence
n!/3!*(n-3)! = 20
n(n-1)(n-2) = 120
(n²-n)(n-2) = 120
n³ - 3n² +2n -120 = 0
Now find n
 
Regards
Arun (askIITians forum expert)